OpenStudy (anonymous):

Will fan & Medal! Admission to the Pie Eating Contest is $1.50 for children and $4 for adults. On a certain day, 2200 people came to watch and $5050 is collected. How many children and adults attended the Contest? Show ALL your work. Let x = number of children Let y = number of adults A. Write 2 equations to model this system: 10 points B. Solve the first equation for x (get the x alone on 1 side): 5 points C. Use Substitution to solve the equations and determine how many adults and how many children attended: 5 points

OpenStudy (anonymous):

x+y=2200 ...(1) 1.5x+4y=5050 \[\frac{ 15 }{ 10 }x+4y=5050\] \[\frac{ 3 }{ 2 }x+4y=5050\] multiply by 2 3 x+8 y=10100...(2) solve (1) and (2)

OpenStudy (anonymous):

that seems like a lot of gibberish....:/

OpenStudy (anonymous):

from (1) x=2200-y put in (2)

OpenStudy (anonymous):

putting in 2...(2)

OpenStudy (anonymous):

3(2200-y)+8y=10100 6600-3y+8y=10100 5y=10100-6600=3500 y=3500/5=700 x=2200-700=1500